Let u(x, t) be a solution of the one- dimensional wave equation and A, B, C and D be any vertices of any parallelogram whose sides are characteristics wave (i.e the line parallel to x+ct=0, x-ct=0)

Then show that u(A)+u(C)=u(B)+u(D)


1
Expert's answer
2021-02-03T14:32:34-0500

A general solution:

"u(x,t)=F(x-ct)+G(x+ct)"

Consider any parallelogram of vertices A, B, C, D with sides parallel to the

characteristics "x=\pm ct+\xi"

Let

"A=(x,t)"

"B=(x+cs, t+s)"

"C=(x+cs-c\tau,t+s+\tau)"

"D=(x-c\tau,t+\tau)"

be the coordinates of the vertices of a characteristic parallelogram, where "s" and "\tau" are

positive parameters.

Then:

"u(A)=F(x-ct)+G(x+ct)"

"u(C)=F(x-2c\tau-ct)+G(x+2cs+ct)"

"u(B)=F(x-ct)+G(x+2cs+ct)"

"u(D)=F(x-2c\tau-ct)+G(x+ct)"

Therefore,

"u(A)+u(C)=u(B)+u(D)"